# Holistic Optimization Framework for FPGA Accelerators

### v4

@@ 3.2. Data-tiling and Padding
To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. In Listing 4, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
@@ 3.3. Fine-grained Parallelism
Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops. Using ISCC (Verdoolaege, 2011), we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is not feasible, the inter-tile loop retains its legal position. For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction. For instance, in Listing 4, the inter-tile loops on lines 3 and 5 iterate over task0 and task1. The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 4, additions take 3 cycles, resulting in $II=3$.
@@ 3.4. Loop Order
Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP (cf. Section 4). As shown in Listing 4, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 3.
@@ 4. NLP Formulation
$II_{l}$ II of the loop $l$
$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
@@
$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation $op$
$DSP$ Number of DSP available for the FPGA used
@@
$SLR$ Number of SLR available for the FPGA used
@@ 4.1. Constraints
We now describe the constraints by using the code of Listings 3, 4 and 5.
@@ 8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 5: Implementation of Task 3 in the 3mm Kernel (from Listing 4). This listing illustrates the structure of Task 3, where the intra-tile computation is fully unrolled to expose fine-grained parallelism. The unrolling enables concurrent execution of loop iterations.
@@ 4.1.1. Data-tiling and Unroll Factor
The intra-tile transformation, as explained in Section 3, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 5, for the array $F$, the loop $j$ (line 7 in Listing 3) has been split into $j0$ (line 14 in Listing 4) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
@@ 4.1.4. Transfer and Reuse
The array $E$, defined on line 24 in Listing 4, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 6 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 4, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
@@ 4.1.6. Array Partitioning
Array D in Listing 5 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
@@ 4.2. Objective Function
The objective function models the dependency graph of the fused task. Independent tasks execute simultaneously, with latency set by the longest task. Dependent tasks overlap partially, with the second task starting once it receives enough data from the first. The latency is the maximum of both tasks’ durations plus the “shift”, the interval before the second task starts. For our 3mm example, the graph is shown in Figure 2, and the corresponding latency is:
$\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
PrometheusT0T1T2T3T4T5
Figure 2. Maximally distributed dataflow graph of the 3mm kernel, where each node represents a task and each edge indicates a data dependency between tasks.
The latency of the intra-task for a statement $s$ is determine similarly to (Pouget et al., 2025a, 2024, b) with Equation 12. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 13.
(12) $Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$
@@
(13) $Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$
For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication. The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
(14) $Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$
And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
@@ 5.2. Communication between the Fused Task
The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
@@ 6.1. Setup
We evaluated our method using kernels from Polybench/C 4.2.1 (Pouchet and Yuki, 2024) with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. Due to Allo’s lack of automatic code generation, we limited our experiments to the subset of PolyBench kernels provided in its artifact evaluation package (all, 2024), as results for other kernels were not available. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow (AMD/Xilinx, 2024). This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.

### v5

@@ 3.1. Dependency Graph Creation
Figure 3 shows the task graph of the 3mm kernel (Listing 5), where each loop body corresponds to a distinct computation task.
T0T1T2T3T4T5
Figure 3. Dataflow graph of the 3mm kernel. Nodes represent computation tasks; edges denote data dependencies.
@@ 3.2. Data-tiling and Padding
To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. In Listing 6, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
@@ 3.3. Fine-grained Parallelism
Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops. Using ISCC (Verdoolaege, 2011), we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is not feasible, the inter-tile loop retains its legal position. For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction. For instance, in Listing 6, the inter-tile loops on lines 3 and 5 iterate over task0 and task1. The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 6, additions take 3 cycles, resulting in $II=3$.
@@ 3.4. Loop Order
Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP (cf. Section 4). As shown in Listing 6, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 4.
@@ 4. NLP Formulation
$I_{l} II$ of the loop $l$
$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without (parpar) and with (redred) dependencies of the statement $s$
@@
$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation opop
DSPDSP Number of DSP available for the FPGA used
@@
SLRSLR Number of SLR available for the FPGA used
@@ 4.1. Constraints
We now describe the constraints by using the code of Listings 4, 6 and 7.
@@ 8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 7: Implementation of Task 3 in the 3mm Kernel (from Listing 6). This listing illustrates the structure of Task 3, where the intra-tile computation is fully unrolled to expose fine-grained parallelism. The unrolling enables concurrent execution of loop iterations.
@@ 4.1.1. Data-tiling and Unroll Factor
The intra-tile transformation, as explained in Section 3, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 7, for the array $F$, the loop $j$ (line 7 in Listing 4) has been split into $j0$ (line 14 in Listing 6) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
@@ 4.1.4. Transfer and Reuse
The array $E$, defined on line 24 in Listing 6, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 6 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 6, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
@@ 4.1.6. Array Partitioning
Array D in Listing 7 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
@@ 4.2. Objective Function
The objective of our framework is to minimize the total latency of the design, which is modeled as a directed acyclic graph (DAG). In this dataflow graph, each node represents a computation task (or fused task), and edges represent data dependencies between tasks.
To compute the overall latency, we define the latency of each task $T\in\mathcal{T}$ as the sum of:
  • the time at which it becomes ready to start (determined by its dependencies)
  • and the duration of the task itself
We denote:
  • $\text{Lat}(T)$: the global latency contribution of task $T$, i.e., the time at which $T$ finishes,
  • $\text{Lat}_{\text{task}}(T)$: the execution time (duration) of task $T$,
  • $\text{pred}(T)$: the set of immediate predecessor tasks of $T$,
  • $\text{shift}_{T_{i},T}$: the number of cycles after which $T$ can start once $T_{i}$ has begun (e.g., due to pipelined data production).
Then the latency of each task is recursively defined as:
(12) $\text{Lat}(T)=\max_{T_{i}\in\text{pred}(T)}\left[\text{Lat}(T_{i})+\text{shift}_{T_{i},T}\right]+\text{Lat}_{\text{task}}(T)$
The overall latency of the design is defined as the latest finish time among all sink tasks (i.e., tasks without successors):
(13) $\text{Lat}_{\text{total}}=\max_{T\in\mathcal{S}}\left[\text{Lat}(T)\right]$
where $\mathcal{S}\subseteq\mathcal{T}$ is the set of sink tasks in the graph.
For the graph in Figure 3, the latency can be expressed step-by-step as:
$\left.\begin{aligned} &\text{Lat}(T_{1})=\text{Lat}(T_{0})+\text{shift}_{T_{0},T_{1}}+\text{Lat}_{\text{task}}(T_{1})\\ &\text{Lat}(T_{3})=\text{Lat}(T_{2})+\text{shift}_{T_{2},T_{3}}+\text{Lat}_{\text{task}}(T_{3})\\ &\text{Lat}(T_{5})=\max\left(\text{Lat}(T_{1})+\text{shift}_{T_{1},T_{5}},\;\text{Lat}(T_{3})+\text{shift}_{T_{3},T_{5}},\;\text{Lat}(T_{4})+\text{shift}_{T_{4},T_{5}}\right)+\text{Lat}_{\text{task}}(T_{5})\end{aligned}\right$.
In the next section, we explain how to compute $\text{Lat}_{\text{task}}(T)$ for each task using intra-task and inter-tile modeling.
4.2.1. Intra-Task and Multi-Level Latency Modeling
Each task $T$ is composed of computation over a tile of data and may be enclosed in multiple inter-tile loop levels (e.g., tiling for cache reuse). At each level, we must account for both computation and communication (i.e., data transfers).
We model the execution as a combination of:
  • Load: reading data from off-chip memory (or previous task) into on-chip buffers,
  • Compute: performing the actual tile computation,
  • Store: writing the result back to off-chip memory (or to a subsequent task).
To enable communication/computation overlap, we use double- or triple-buffering depending on the number of streams. At each inter-tile level $n$, the total latency is computed as the maximum of the three components, accounting for pipeline shifts and buffer reuse.
Level-Based Latency Recursion.
Let:
  • $\text{Lat}_{n+1}$: latency of executing the inner level $n+1$,
  • $f_{a,n}$: number of bytes transferred for array $a$ at level $n$,
  • $BS_{a}$: memory bandwidth for array $a$,
  • $\alpha$: overlap factor (1 = load or store only, 2 = both).
Then, the latency at level $n$ is:
(14) $\text{Lat}_{n}=\max\left(\text{Lat}_{n+1},\frac{f_{a,n}}{BS_{a}}\right)+\text{Lat}_{n+1}+\alpha\cdot\frac{f_{a,n}}{BS_{a}}$
This formula captures both the initialization overhead and the overlap between communication and computation.
Base Case: Intra-Task Latency.
The base latency $\text{Lat}_{n+1}$ is the latency of the innermost computation tile, computed as in (Pouget et al., 2025a, b).
(15) $\text{Lat}_{\text{intra}}=IL_{\text{par}}+IL_{\text{seq}}\cdot\log_{2}\left(\prod_{l\in\mathcal{L}^{\text{red}}_{s}}TC^{l}_{\text{intra}}\right)$
If the computation is invoked across multiple tiles in a pipelined loop, the full latency becomes:
@@ 4.2.1. Intra-Task and Multi-Level Latency Modeling
(16) $\text{Lat}_{\text{task}}=\text{Lat}_{\text{intra}}+II\cdot\left(\prod_{l\in\mathcal{L}^{\text{red}}_{s}}TC^{l}_{\text{inter}}-1\right)$
Final Task Latency.
The full latency of a task $T$, denoted $\text{Lat}_{\text{task}}(T)$, is the latency at the outermost loop level $n=0$, computed recursively using Equation 14, where the base is given by Equation 16.
@@ 5.1. Communication with Off-Chip Memory
Listing 8 shows the load and read functions, which first transfer data from off-chip memory to an on-chip FIFO, and then read from the FIFO into the local buffer allocated for Task 1.
1void load_vA_for_task1(hls::stream<float16> &fifo_A_from_off_chip_to_S1,
2 float16 vA[2340]) {
3#pragma HLS inline off
4 for (int i = 0; i < 2340; i++) {
5#pragma HLS pipeline II = 1
6 fifo_A_from_off_chip_to_S1.write(vA[i]);
7 }
8}
9
10void read_A_FT0(float A[10][204],
11 hls::stream<float16> &fifo_A_from_off_chip_to_S1, int i0) {
12#pragma HLS inline off
13 if (i0 >= 18) {
14 return;
15 }
16 for (int d0 = 0; d0 < 10; d0++) {
17 for (int d1 = 0; d1 < 204; d1 += 16) {
18#pragma HLS pipeline II = 1
19 float16 tmp_fifo = fifo_A_from_off_chip_to_S1.read();
20 for (int j = 0; j < 16; j++){
21#pragma HLS unroll
22 if (d1 + j < 204)
23 A[d0][d1 + j] = tmp_fifo[j];
24 }
25 }
26 }
27}
Listing 8: Memory transfer for array A in the 3mm example (Listing 4). The data is first loaded from off-chip memory into a FIFO, then read from the FIFO to populate the on-chip buffer
@@ 5.2. Communication between the Fused Task
The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap. Listing 9 illustrates an example where the write function transfers data produced by Tasks 0 and 1 into a FIFO, which is then consumed by Task 5. As shown, the on-chip buffers used by Tasks 0 and 1 differ from those used by Task 5, allowing greater flexibility in choosing tile sizes for each task independently.
1void write_E_FT0(float E[10][32],
2 hls::stream<float16> &fifo_E_from_task1_to_task5, int j0,
3 int i0) {
4#pragma HLS inline off
5 if (j0 < 0 $|$ i0 < 0) {
6 return;
7 }
8 for (int d0 = 0; d0 < 10; d0++) {
9 for (int d1 = 0; d1 < 32; d1 += 16) {
10#pragma HLS pipeline II = 1
11 float16 tmp_fifo;
12 for (j = 0; j < 16; j++){
13#pragma HLS unroll
14 tmp_fifo[j] = E[d0][d1 + j];
15 }
16 fifo_E_from_task1_to_task5.write(tmp_fifo);
17 }
18 }
19}
20
21void read_E_FT2(float E[180][192],
22 hls::stream<float16> &fifo_E_from_task1_to_task5, int i0,
23 int j0) {
24#pragma HLS inline off
25 if (j0 > 0 $|$ i0 >= 18) {
26 return;
27 }
28 for (int d1_0 = 0; d1_0 < 6; d1_0++) {
29 for (int d0 = 0; d0 < 10; d0++) {
30 for (int d1_1 = 0; d1_1 < 32; d1_1 += 16) {
31 int d1 = d1_0 * 32 + d1_1;
32 float16 tmp_fifo = fifo_E_from_task1_to_task5.read();
33 for (j = 0; j < 16; j++){
34#pragma HLS unroll
35 if (d1 + j < 192)
36 E[d0 + i0 * 10][d1 + 0 + j] = tmp_fifo[j];
37 }
38 }
39 }
40 }
41}
Listing 9: FIFO-Based Inter-Task Communication for Intermediate Matrix E: The first function writes data from the local buffer E into a FIFO, while the second reads from the FIFO to populate the next computation stage’s on-chip buffer.
@@ 6.1. Setup
We evaluated our method using kernels from Polybench/C 4.2.1 (Pouchet and Yuki, 2024) with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. Due to Allo’s lack of automatic code generation, we limited our experiments to the subset of PolyBench kernels provided in its artifact evaluation package (all, 2024), as results for other kernels were not available. We include a comparison with Sisyphus using the n-madd kernels, where $n$ denotes the number of matrix additions. In the case of 2-madd, the result of the first addition is used as the input to the second. For 3-madd, the results of the first two additions are both used as inputs to the final addition.
Table 5 lists the benchmark kernels used in our evaluation. For each kernel, we report the computational complexity (in terms of the number of operations) and the memory complexity (in terms of footprint of the input/output data). Additionally, we include the data reuse order, which reflects the approximate number of times one data element is reused across different computations within the kernel. Kernels with reuse on the order of $\mathcal{O}({N})$, where $N$ denotes the problem size (e.g., the number of rows or columns in an $N\times N$ input matrix), are typically considered compute-bound as their theoretical arithmetic intensity (assuming perfect reuse on-chip) is $\mathcal{O}({N})$. In contrast, kernels with reuse on the order of $\mathcal{O}({1})$ are generally classified as memory-bound. In practice, compute-bound kernels require very careful bufferization on-chip to minimize off-chip communications and achieve the theoretical arithmetic intensity. The final column (Communication Between Tasks) represents the number of data elements that are transferred between tasks in the dataflow design, excluding any initial input loading or setup overhead.
Benchmark Description Ops      Complexity Mem      Complexity Reuse Comm. Between Tasks
bicg BiCG sub-kernel of BiCGStab solver $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ 0
madd Matrix add. (C = A + B) $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ 0
mvt Matrix Vector product and Transpose $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ 0
atax Matrix transpose and vector mult. $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ NN
gesummv Scalar, vector and matrix mult. $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ $2N$
2-madd 2 Matrix add. (D = (A + B) + C) $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ $N^{2}$
[…]
gemver Vector mult. and matrix add. $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ $2N^{2}+2N$
2mm 2 Matrix Mult. ($\alpha ABC+\beta D$) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
gemm Matrix-multiply ($C=\alpha AB+\beta C$) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
syr2k Symmetric rank-2k update $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
syrk Symmetric rank-k update $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
trmm Triangular matrix-mult. $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
[…]
symm Symmetric matrix-mult. $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $2N^{2}$
Table 5. Benchmark Kernels Used for Evaluation: Computational Complexity, Memory Requirements, Data Reuse, and Inter-Task Communication Analysis
NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow (AMD/Xilinx, 2024). This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.